# DFA-GNN: Forward Learning of Graph Neural Networks by Direct Feedback Alignment

### v1

@@ 4.2 Pseudo Error Generator by Spreading Residual Errors
In the computation of gradient (Eq. 6), the labels of all nodes are required, but not all nodes are labeled in the semi-supervised learning task. To address this issue, we introduce a mechanism to generate errors, assigning a pseudo error to each unlabeled node. Inspired by the ingenious work of [18] which makes correction for the prediction of simple base predictor, we adopt the simlilar approach for the generation of pseudo error for each unlabeled node. The underlying principle is the expectation that errors in initial predictions are likely to propagate along the graph edges. That is, an error at a given node $v$ suggests a higher likelihood of similar errors to its the neighbor nodes. This concept of error propagation across the graph is supported by previous studies [19]. Our approach draws inspiration from the strategy of residual propagation used in node regression tasks, and more broadly, from the frameworks of generalized least squares and correlated error models [20].
@@
$\textbf{Z}^{*}=\arg\min_{\textbf{Z}\in\mathbb{R}^{n\times c}}tr(\textbf{Z}^{\text{T}}(\textbf{I}-\textbf{S})\textbf{W})+\mu\|\textbf{Z}-\textbf{E}\|_{F}^{2},$ (7)
where $tr(\cdot)$ denotes the trace of a matrix. The first term enhances the smoothness of the error estimation throughout the graph, while the second term ensures that the final solution stays consistent with the initial error estimate E. Following the optimization methodology in [21], the solution to Eq. 7 can be obtained through iterative processing
@@ A.4 Datasets Statistics
We evaluate our method on 10 benchmark datasets across domains: citation networks (Cora, CiteSeer, PubMed) [30, 31], Amazon co-purchase graph (Photo, Computer) [32], Wikipedia graphs (Chameleon, Squirrel) [33], actor co-occurrence graph (Actor) [34] and webpage graphs from WebKB (Texas, Cornell) [34]. The datasets adopted are representative which describe diverse real-life scenarios. Some of them are highly homophilic while others are heterophilic. Note that for Chameleon and Squirrel, we use the filtered version from [35], as the original version from [33] may contain duplicated nodes. The detailed statistics of the datasets are summarized in Tab. 4. We compute the node homophily for each dataset using the method proposed by [34], referred to as the homophily value.
@@
  • Cora, CiteSeer and PubMed [30] are three classic homophilic citation networks. In these networks, nodes correspond to academic papers, and edges signify the citation links between papers. The node features are derived from bag-of-word representations of the papers, and the labels categorize each paper into specific research topics.
  • Computer and Photo [32] are segments of the Amazon co-purchase graph, where nodes represent goods, edges indicate that two goods are frequently bought together, node features are bag-of-words encoded product reviews, and class labels are given by the product category.
  • Chameleon and Squirrel [33] are two heterophilic networks derived from Wikipedia. In these networks, nodes represent Wikipedia web pages, and edges correspond to hyperlinks between these pages. The features are comprised of informative nouns extracted from the Wikipedia content, while the labels reflect the average traffic of each web page.
  • Actor [34] is a heterophilic actor co-occurrence network where nodes represent actors, and edges signify that two actors have appeared together in the same movie. The features are derived from keywords found on the actors’ Wikipedia pages, while the labels consist of significant words associated with each actor.
  • Cornell and Texas [34] are heterophilic networks from the WebKB1 project representing computer science departments at three universities. Nodes are departmental web pages, edges represent hyperlinks, features are derived using bag-of-words, and labels categorize page types. These networks illustrate heterophilic connections where linked pages often differ in type.
@@ A.5 Experimental Settings
The codes of DFA-GNN are based on the GNNs in the PyTorch version by Deep Graph Library (DGL) [36]. To generate pseudo errors, we search the optimal $\alpha$ in Eq. 8 within {0.001, 0.01, 0.1, 0.3, 0.5, 0.7, 0.9}, the iteration epoch for Eq. 8 within {5, 10, 30, 50, 100, 150, 200}, $\epsilon$ in Eq. 9 within {0.3, 0.5, 0.7, 0.9}. For the training of our method, we search the learning rate within {0.001, 0.01, 0.1} and weight decay within {0.0005, 0}. All the experiments are run on AMD EPYC 7542 32-Core Processor with Nvidia GeForce RTX 3090. We list the hyper-parameter values used in our model in Tab. 5.
@@ A.6 Time Comparison
A.7 Codes of baseline models
We implemented all baseline models using the official code and leveraged libraries such as PyTorch Geometric [37], DGL [36], and Deeprobust [24]. The URLs of all the baseline approaches we used are listed in Tab. 7. Our code will be made public upon the acceptance of this paper.
Table 7: Codes of baseline approaches.
URL
FF https://github.com/mpezeshki/pytorch_forward_forward
FORWARDGNN https://github.com/facebookresearch/forwardgnn
PEPITA https://github.com/GiorgiaD
CaFo https://github.com/Graph-ZKY/CaFo
GCN https://github.com/tkipf/gcn
SGC https://github.com/Tiiiger/SGC
APPNP https://github.com/benedekrozemberczki/APPNP
GraphSage https://github.com/williamleif/GraphSAGE
GAT https://github.com/PetarV-/GAT
ChebNet https://github.com/dsgiitr/graph_nets

### v2

@@ 4.2 Pseudo Error Generator by Spreading Residual Errors
In the computation of gradient (Eq. 6), the labels of all nodes are required, but not all nodes are labeled in the semi-supervised learning task. To address this issue, we introduce a mechanism to generate errors, assigning a pseudo error to each unlabeled node. The underlying principle is the expectation that errors in initial predictions are likely to propagate along the graph edges. That is, an error at a given node $v$ suggests a higher likelihood of similar errors to its the neighbor nodes. This concept of error propagation across the graph is supported by previous studies [18]. Our approach draws inspiration from the strategy of residual propagation used in node regression tasks, and more broadly, from the frameworks of generalized least squares and correlated error models [19].
@@
$\textbf{Z}^{*}=\arg\min_{\textbf{Z}\in\mathbb{R}^{n\times c}}tr(\textbf{Z}^{\text{T}}(\textbf{I}-\textbf{S})\textbf{Z})+\mu\|\textbf{Z}-\textbf{E}\|_{F}^{2},$ (7)
where $tr(\cdot)$ denotes the trace of a matrix. The first term enhances the smoothness of the error estimation throughout the graph, while the second term ensures that the final solution stays consistent with the initial error estimate E. Following the optimization methodology in [20] and [21], the solution to Eq. 7 can be obtained through iterative processing
@@ 5.2 Ablation Study
5.3 Visualization of Convergence
To better illustrate the training convergence of DFA-GNN, we plot the training and validation accuracy of the proposed DFA-GNN and BP over training epochs on three datasets, as shown in Fig. 3. In general, our method shows similar convergence to BP. For both BP and our method, the convergence of validation accuracy occurs much earlier than that of training accuracy due to overfitting. The validation convergent epoch of DFA-GNN is nearly the same as BP on Cora and CiteSeer (around 100 epochs), while it is 100 epochs later than BP on PubMed (around 200 epochs). Our method achieves better validation accuracy on all these datasets and suffers less from overfitting compared with BP. In terms of training accuracy, the convergence of our method is slightly slower than BP because the update direction of our method is not exactly opposite to the gradient direction but maintains a small angle. Since our method considers both the errors of labeled nodes and pseudo errors of unlabeled nodes as supervision information, which is different from BP that only uses the loss of labeled nodes for supervision, the convergent value of training accuracy for our method is slightly lower than BP. However, this does not affect our method achieving better validation results.
(a) Cora
(b) CiteSeer
(c) PubMed
Figure 3: Visualization of the convergence of BP and our method on Cora, CiteSeer, and PubMed.
@@ 5.4 Robustness Analysis
Table 3: Results on three large datasets. OOM denotes out of memory.
FF+LA FF+VN PEPITA CaFo+MSE CaFo+CE SF BP Ours
Flickr 6.09 42.40 49.28 50.02 49.69 46.47 50.79 49.80
Reddit 12.44 OOM OOM 88.15 91.55 94.38 94.34 94.49
ogbn-arxiv 56.38 19.84 35.16 53.51 60.57 66.54 68.78 67.83
@@
5.5 Scalability on Large Datasets
Our method is well-suited for large datasets. When the graph scale is large, we can use edge indices instead of an adjacency matrix to store the graph. For forward propagation (Eq. 1), the complexity of neighbor aggregation can be reduced from $\mathcal{O}(n^{2}d)$ to $\mathcal{O}(|\mathcal{E}|d)$, where $|\mathcal{E}|$ denotes the number of edges. For direct feedback alignment (Eq. 10), as $(\textbf{S}^{\text{T}})^{k}\hat{\textbf{E}}$ is exactly the aggregation of errors for $k$ times, the time and space complexity can be reduced to $\mathcal{O}(kc|\mathcal{E}|)$, without the need to calculate the $k$-th power of the adjacency matrix. Similarly, complexity reduction can also be achieved in the node filtering process. The experimental results on the Flickr [25], Reddit [26], and ogbn-arxiv [27] datasets, as presented in Tab. 3, demonstrate that our method is effective on large-scale datasets, delivering strong performance. Our method achieves results comparable to BP while surpassing other non-BP methods, all with a small memory footprint (2043 MiB for Flickr, 11675 MiB for Reddit, and 2277 MiB for ogbn-arxiv). This indicates that our method not only scales well to large graphs but also maintains efficiency in terms of space usage.
@@ A.4 Datasets Statistics
We evaluate our method on 10 benchmark datasets across domains: citation networks (Cora, CiteSeer, PubMed) [sen2008collective, yang2016revisiting], Amazon co-purchase graph (Photo, Computer) [mcauley2015image], Wikipedia graphs (Chameleon, Squirrel) [rozemberczki2021multi], actor co-occurrence graph (Actor) [pei2019geom] and webpage graphs from WebKB (Texas, Cornell) [pei2019geom]. The datasets adopted are representative which describe diverse real-life scenarios. Some of them are highly homophilic while others are heterophilic. Note that for Chameleon and Squirrel, we use the filtered version from [platonov2022critical], as the original version from [rozemberczki2021multi] may contain duplicated nodes. The detailed statistics of the datasets are summarized in Tab. 5. We compute the node homophily for each dataset using the method proposed by [pei2019geom], referred to as the homophily value.
@@
  • Cora, CiteSeer and PubMed [sen2008collective] are three classic homophilic citation networks. In these networks, nodes correspond to academic papers, and edges signify the citation links between papers. The node features are derived from bag-of-word representations of the papers, and the labels categorize each paper into specific research topics.
  • Computer and Photo [mcauley2015image] are segments of the Amazon co-purchase graph, where nodes represent goods, edges indicate that two goods are frequently bought together, node features are bag-of-words encoded product reviews, and class labels are given by the product category.
  • Chameleon and Squirrel [rozemberczki2021multi] are two heterophilic networks derived from Wikipedia. In these networks, nodes represent Wikipedia web pages, and edges correspond to hyperlinks between these pages. The features are comprised of informative nouns extracted from the Wikipedia content, while the labels reflect the average traffic of each web page.
  • Actor [pei2019geom] is a heterophilic actor co-occurrence network where nodes represent actors, and edges signify that two actors have appeared together in the same movie. The features are derived from keywords found on the actors’ Wikipedia pages, while the labels consist of significant words associated with each actor.
  • Cornell and Texas [pei2019geom] are heterophilic networks from the WebKB1 project representing computer science departments at three universities. Nodes are departmental web pages, edges represent hyperlinks, features are derived using bag-of-words, and labels categorize page types. These networks illustrate heterophilic connections where linked pages often differ in type.
@@ A.5 Experimental Settings
The codes of DFA-GNN are based on the GNNs in the PyTorch version by Deep Graph Library (DGL) [wang2019deep]. To generate pseudo errors, we search the optimal $\alpha$ in Eq. 8 within {0.001, 0.01, 0.1, 0.3, 0.5, 0.7, 0.9}, the iteration epoch for Eq. 8 within {5, 10, 30, 50, 100, 150, 200}, $\epsilon$ in Eq. 9 within {0.3, 0.5, 0.7, 0.9}. For the training of our method, we search the learning rate within {0.001, 0.01, 0.1} and weight decay within {0.0005, 0}. All the experiments are run on AMD EPYC 7542 32-Core Processor with Nvidia GeForce RTX 3090. We list the hyper-parameter values used in our model in Tab. 6.
@@ A.6 Time Comparison
A.7 Comparison of BP and DFA with Pseudo-Error Generation
Although the pseudo-error generation process is essential for our method, it is also an optional choice for backpropagation. We integrate this component into BP, and the experimental results from Tab. 8 show that although this component contributes to DFA in our method, it does not positively enhance BP overall. Even with pseudo-error generation, BP cannot outperform our method. This observation indicates that direct feedback of errors may benefit more from pseudo-errors rather than the layer-by-layer backward pass.
Table 8: Results of BP and DFA with pseudo-error generation spreading: mean accuracy (%)$\pm$95confidence interval.
Cora CiteSeer PubMed Photo Computer Actor Chameleon Squirrel
BP
86.04
$\pm$0.62
78.20
$\pm$0.57
85.24
$\pm$0.28
93.03
$\pm$0.59
89.48
$\pm$0.37
31.94
$\pm$0.88
41.28
$\pm$2.29
37.81
$\pm$0.71
BP+EG
87.41
$\pm$0.80
(1.37$\uparrow$)
80.24
$\pm$1.11
(2.04$\uparrow$)
84.74
$\pm$0.41
(0.50$\downarrow$)
91.95
$\pm$0.41
(1.08$\downarrow$)
87.78
$\pm$0.51
(1.70$\downarrow$)
31.53
$\pm$1.91
(0.41$\downarrow$)
38.71
$\pm$2.23
(2.57$\downarrow$)
35.78
$\pm$1.52
(2.03$\downarrow$)
DFA
83.02
$\pm$1.36
78.17
$\pm$0.71
82.92
$\pm$0.42
91.75
$\pm$0.48
84.02
$\pm$1.54
30.72
$\pm$1.72
39.09
$\pm$1.17
34.61
$\pm$1.01
DFA+EG (ours)
87.72
$\pm$1.63
(4.70$\uparrow$)
80.49
$\pm$0.41
(2.32$\uparrow$)
86.28
$\pm$0.67
(3.36$\uparrow$)
93.04
$\pm$0.31
(1.29$\uparrow$)
86.72
$\pm$0.68
(2.70$\uparrow$)
34.07
$\pm$0.75
(3.35$\uparrow$)
41.19
$\pm$1.56
(2.10$\uparrow$)
38.17
$\pm$2.21
(3.56$\uparrow$)
A.8 DFA-GNN with Alternative Activation Functions
We conduct experiments for our method with four different activation functions (i.e., Sigmoid, Tanh, ELU and LeakyReLU) as shown in Tab. 9. The results demonstrate our method is well integrated with different activation functions and derives consistently good results.
Table 9: Results of our method with different activation functions.
ours+Sigmoid ours+Tanh ours+ReLU ours+ELU
ours+LeakyReLU
(slope=0.2)
Cora 87.68$\pm$1.75 88.17$\pm$1.66 87.72$\pm$1.63 87.93$\pm$1.75 87.64$\pm$1.65
CiteSeer 79.94$\pm$0.74 80.01$\pm$0.75 80.49$\pm$0.41 81.11$\pm$0.86 80.43$\pm$0.75
PubMed 84.57$\pm$0.32 85.57$\pm$0.40 86.28$\pm$0.67 85.53$\pm$0.48 84.99$\pm$0.39